# Gatsby without the ‘E’: Creating Lipograms with LLMs

### v1

@@ 1 Introduction
The lipogram is a fascinating literary form, requiring the writer to omit all occurrences of a particular letter from the text. Perhaps the best known lipogram is Gadsby by Ernest Vincent Wright (Wright, 1939), which systematically excludes the letter ‘e’ throughout his novel of over 50,000 words. Crafting such a lipogram is challenging, because ‘e’ is the most common letter in English, representing roughly 12% of the letters appearing in conventional English-language text.
Lipograms are interesting as a highly specific literary form, one that provides insight into the flexibility and malleability of language under strong constraints. Modern large language models (LLMs) excel in their ability to generate text under strong stylistic constraints, such as replicating the style of Shakespeare or Dr. Seuss. This makes lipogram generation an excellent stress test for LLMs, which must navigate these literary constraints without sacrificing fluency or meaning.
@@
Our particular challenge is to translate F. Scott Fitzgerald’s classic American novel “The Great Gatsby” Fitzgerald (1925) into an ‘e’-less lipogram. Gatsby was selected as our target because it is a short (47,094 word) novel familiar to a broad readership, and with the recent expiration of copyright protection now in the public domain. Finally, the name Gatsby resonates well with Gadsby, the lipogram novel which is its rough contemporary.
Our contributions in this paper include:
  • Methods for Lipogram Translation: We present a systematic study of algorithms for creating lipograms, from baseline methods such as synonym replacement to sophisticated generative models enhanced using (i) beam search, (ii) name entity analysis, (iii) training on paraphrased texts, (iv) optimization by sequence trimming and (v) multi-selection. We present the results of an ablation study that assesses the relative effectiveness for many of these techniques.
  • Lipogram Evaluation: We measure the quality of generated lipogram translations across a variety of metrics, including readability, grammatical correctness, and semantic similarity to the original source text. Beyond powering our feature ablation study, we use this evaluation suite to compare the performance of three different LLMs (T5, Llama2, and Llama3), demonstrating that that Llama3 produced the most faithful as well as the most readable constrained translations. These metrics and the data resources we will provide open up a new challenge for future researchers to produce better lipogram translations.
Our project not only demonstrates the feasibility of creating high-quality lipograms using advanced language models, but also provides valuable insights into the complexities and potential solutions in constrained text generation. We provide the first page of our lipogram translation of Gatsby for inspection in Section B of the appendix, along with the original source text. Upon acceptance of this paper we will release our full lipogram translation and all associated data/code.
This paper is organized as follows. We survey previous work on lipograms and constrained text generation in Section 2. Our lipogram generation and evaluation methods are presented in Sections 3 and 4, respectively. We present our results on the malleability of English in Section 5, with suggestions for future work.
2 Previous Work
We review the literary history of lipograms and NLP research on constrained text generation in the following subsections.
2.1 Literary and Popular Discussion
  • La Disparition (A Void) (Perec, 1969) by Georges Perec is a French novel inspired by "Gadsby" that also excludes the letter ‘e’. It was later translated into English by Gilbert Adair while maintaining this constraint. The narrative follows the mysterious disappearance of a certain Anton Vowl, weaving a complex plot without using ‘e’. Perec followed up with his novel “Les Revenentes” Perec (1972) which uses only words with the vowel ‘e’, showcasing his exploration of constrained writing beyond lipograms.
  • Eunoia (Bök, 2001) by Christian Bök is a univocalic lipogram, where each chapter uses only one vowel. It explores the creative possibilities of writing with severe constraints, producing distinct and innovative prose.
The concept of constrained writing has garnered significant attention in both literary circles and popular media. The Oulipo (Ouvroir de littérature potentielle) group, co-founded by Raymond Queneau and François Le Lionnais, was instrumental in exploring the potential of literary constraints. Their works, including Perec’s La Disparition (Perec, 1969) and Calvino’s If on a winter’s night a traveler (Calvino and Weaver, 1993), exemplify the creative possibilities that constraints can unlock in literature Motte (1986).
Authors have experimented with various other constraints in their writing beyond omiting the letter ‘e’, Alphabetical constraints in writing began with Tryphiodorus’ adaptation of the Odyssey, written between the 2nd and 4th centuries, omits successive letters of the Greek alphabet in each book, such as ‘alpha’ in the first book, ’beta’ in the second, and so on. (Wikipedia, 2024) This is analogous to the lipogram translation task we set out in this paper.
A related challenge is that of producing palindromic text. There exists a small genre of works such as “Dr. Awkward & Olson in Oslo” Levine (1986); Puder (2016) where the text reads the same backward as forward.
2.2 Technical Generation of Lipograms
Recent advancements in natural language processing (NLP) have led to innovative methods for generating constrained text. Roush et al. (2023) introduces a technique for modifying the output of language models by applying filter functions to their vocabularies before generating text. This method, implemented in the Constrained Text Generation Studio (CTGS) and the Gadsby web app, allows users to generate text with various constraints, including banning specific letters and enforcing word patterns. The CTGS tool demonstrates the robustness of language models to constraints, and shows that filtering tokens before generation leads to better performance and eliminates constraint errors compared to fine-tuning alone.
Kumar et al. (2021) presented a method for Controlled Text Generation by treating it as an optimization problem. They used gradient descent to maximize both the log probability of the language model and the constraint objectives, focusing exclusively on sequence-level constraints.
Zhang et al. (2020) tackled the challenge of hard-constraint text generation with an innovative approach. They have introduced POINTER, an insertion-based transformer model designed for hard-constrained text generation. Unlike BERT and GPT-2, POINTER can generate text under specified lexical constraints, like ensuring that certain words appear in a specific order by progressively inserting tokens. This non-autoregressive method allows for parallel token insertion, achieving logarithmic time complexity during inference.
Liu et al. (2021) addressed constrained text generation in the CommonGen task. The researchers propose reinforcement learning with a comprehensive score as the reward, and design a guided decoding method at word, fragment, and sentence levels. Experiments show this approach significantly improves concept coverage and outperforms existing models. However, their approach requires extensive training and isn’t effective for handling strict lexical or phonetic constraints.
3 Generation Approaches
In this section, we discuss our methods for translation under lipogram constraints. We will compare these methods against two simple baselines:
  • Removal of All ‘E’s: Our initial approach simply removes all instances of the letter ‘e’ from the source text. Although straightforward, such an approach creates numerous spelling errors and a significant loss of readability.
  • Synonym Replacement: Here we lemmatized the words and replaced them as necessary with synonyms from Wordnet (Miller, 1995) that do not contain the letter ‘e’. When no suitable synonym was available, the original word was retained after deleting any ‘e’s it contained.
3.1 Prompt Engineering
Initially, our prompt for the language models (LLMs) was ‘‘Rewrite the text without using the letter e." However, this approach led the models to work on a character level, often deleting ‘e’s from words instead of finding suitable replacements that excluded the letter ‘e’. Through extensive trial and error, we refined our approach.
T5 model prompt:
The prompt used for the T5 model was the straightforward: ‘‘paraphrase: {original paragraph}:". The T5 model is generally used for text-to-text tasks, thus the simplicity of the prompt ensures that the model focuses directly on rephrasing the paragraph without unnecessary complexity, which can lead to hallucinations.
Llama model prompts:
For the Llama models, the prompts were designed to provide explicit instructions while maintaining a formal structure. Our Llama prompt was:
Rewrite the following paragraph without using any words containing the letter ‘e’. Preserve the original meaning as much as possible.
Original Paragraph: {paragraph}
Paraphrased Version (without ‘e’):
This shorter prompt reduces the cognitive load on the model, helping to minimize hallucinations and errors. By reducing the instructions, the prompt provides enough guidance to keep the model on track without increasing the prompt length too much. This refinement leads to more consistent output, with fewer instances of deviation from the given constraints, and thus better results.
3.2 Named Entity Analysis
Specific character and place names must remain invariant over the course of a novel, which is not necessarily the case over multiple independent paragraph translations. LLM generated names are often inconsistent; for example, the Gatsby character named “Miss Baker” might suddenly become “Miss Bakar” or “Miss Baking” at different points in the text.
We used named entity recognition (NER) to identify and replace names containing ‘e’ with fixed alternatives, ensuring uniformity throughout the text. We usedliding window technique with a length of 25 words to resolve pronouns. This approach involved identifying the best antecedent for each pronoun within the context, replacing pronouns with their corresponding entities to maintain coherence, using the Stanford NLPs coreference resolution tool Clark and Manning (2016).
3.3 Fine tuning for Paraphrasing
To enhance performance on the paraphrasing task, we fine-tuned our model using the PAWS (Zhang et al., 2019) dataset, which provides pairs of original and paraphrased texts. For fine-tuning the Llama2 model, we utilized QLoRA Dettmers et al. (2024) with 4-bit quantization and the SFTTrainer framework. The quantization helped in making it possible to run on smaller GPUs without sacrificing performance. For T5, the loss function was modified to include penalties for generating words with the letter ‘e’ (LossE) and for maintaining high semantic similarity to the original text (LossSim):
Total $\text{Total Loss}=\lambda_{1}\times\text{Loss}_{\text{E}}+\lambda_{2}\times\text{Loss}_{\text{Sim}}$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the trade-off between avoiding ‘e’ and preserving meaning.
3.4 Constrained Beam Search
To generate text that strictly avoids the letter ‘e’, we employed a constrained beam search approach during decoding. Beam search allows the exploration of multiple potential sequences simultaneously, increasing the likelihood of an optimal output.
We explicitly filtered out tokens containing the letter ‘e’ by including them in the bad_words_ids list, ensuring that no token with the letter ‘e’ could be part of the generated sequence. Key parameters were set to guide the beam search process: the minimum and maximum lengths of the generated sequences were defined to control the output size, and the temperature set to 0.90 to manage the randomness of the sampling process. We specified the number of beams to consider during decoding, set to 20 to widen the search space. Additionally, we configured a no-repeat n-gram size of 3 to avoid repetitive patterns in the output.
3.5 Multiselection
Our model produced several different candidate translations (3 for T5 and Llama2, 10 for Llama3) and then we selected the one with the highest cosine similarity to the original text. This step ensured that the generated text was not only compliant with the constraint but also closely aligned with the original in terms of content and context, minimizing the risk of producing random or irrelevant text.
3.6 LLM Selection
We utilized both encoder-decoder (T5-small) and decoder-only (Llama2-7b-chat-hf and Llama3-8b-Instruct) models in our work. The T5 model employed constrained beam search, named entity recognition (NER), and pronoun resolution, while Llama2 and Llama3 incorporated QLoRA quantization and constrained decoding. For each model, we passed the prompt to the fine-tuned model and selected the best output using beam search. To ensure similar input and output lengths, we set the minimum length to 50% of the input length and the maximum length to 150% of the input length. This approach allowed us to generate e-less text that maintained high semantic similarity to the original content.
3.7 Post Processing
In our post-processing stage, we cleaned and refined the generated text to ensure it was error-free and readable. First, we improved the lipogram text by truncating irrelevant parts from the response: overlong generation (especially with T5) reflected dissatisfying hallucination by the LLM. We progressively deleted the suffix so as to maximize the semantic similarity with the original text. We unified the punctuation of the text by replacing double backticks and multiple single quotes with standard double quotes and fixed spaces around quotes. We adjusted comma spacing, removed long lists of terms, retained only essential punctuation marks, and eliminated empty paragraphs to keep the text flowing smoothly.
We then used LanguageTool to detect and correct any grammatical errors, and the Proselint library to polish the writing style and adjust punctuation spacing for better readability. This post-processing ensured that the final text was not only free of the letter ‘e’ but also relatively clear and coherent.
@@ 3.7 Post Processing
Model /Score Similarity OOV (%) E-Score Grammar Mistakes (%) Readability
Original 1.00 35.56 37.54 1.19 -3.10
E-Delete 0.61 51.32 0.00 29.26 -2.21
Synonym 0.67 40.21 0.00 24.85 -2.78
T5 0.69 35.27 0.00 10.39 -3.15
Llama2 0.65 37.41 0.00 16.40 -3.75
Llama3 0.71 36.37 0.00 1.08 -3.96
@@
Model /Score Similarity OOV (%) E-Score Grammar Mistakes (%) Readability
Complete Model 0.69 35.04 0.00 0.88 -4.00
Model–Finetuning 0.30 63.17 0.00 26.47 -2.21
Model–(Multiselection, BeamS) 0.68 35.74 0.00 0.93 -3.62
Model–Post Processing 0.70 36.53 0.00 2.37 -4.00
Table 3: Ablation studies on our Llama3-based lipogram translation model.
@@ 4 Evaluation Methods
We employed several evaluation methods to assess the quality and effectiveness of our generated text. These aimed to measure various aspects of the text, including semantic similarity, readability, grammatical accuracy, hallucination, and constraint adherence. Table 2 reports each of these metrics for texts generated by each of our models, compared to the original Gatsby.
4.1 Constraint Violations (Presence of ‘E’)
We first evaluated adherence to our primary constraint of excluding the letter ‘e’. This involved calculating the percentage of violations:
[…]
where $W_{e}$ is the number of words having the letter ‘e’ in the generated text, and $W_{t}$ is the total number of words in the generated text.
All of our methods evaluated here succeed in eliminating 100% of the usage of letter ‘e’.
@@ 4.1 Constraint Violations (Presence of ‘E’)
Figure 1: Cosine similarity distributions for lipogram translations. Better models right-shift distributions.
4.2 Cosine Similarity of Embeddings
@@
We used cosine similarity to measure the semantic equivalence between the original source and our generated lipogram translation. This method computes the dot product of embeddings of both texts to quantify how closely they align in terms of semantics. For embeddings, we utilized the the SentenceTransformer model "paraphrase-MiniLM-L6-v2" model to encode semantics. (Reimers and Gurevych, 2019)
Figure 1 presents a histogram of the paragraph similarity over the following generation models: E-delete, Synonym replacement, T5, Llama2, and Llama3 models. Models producing right-shifted peaks indicate better performance, with Llama3 and T5 generating paragraphs more similar to the original, although there remains a sizable gap with the original Gatsby text.
4.3 Out of Vocabulary Score
@@
Lipogram construction algorithms are naturally tempted to delete all occurrences of forbidden symbols, whether they leave a true vocabulary word behind or not (i.e. bhind). To measure this we calculated an out of vocabulary (OOV) score, based on the percentage of words in the generated text that were not found in standard dictionaries. A higher OOV score suggests a greater degree of ill-formed content. We note that 35.56% of the words in original Gatsby qualify as OOV, largely reflecting proper nouns and obsolete terms.
Figure 2 plots the per-paragraph OOV frequency for each model. The E-delete method performs the worst, with over half the words OOV (and many non-OOV terms matching unrelated words). T5 best matched the source distribution, suggesting better vocabulary handling, while the Synonym and Llama2 models also perform well. Llama3 performs significantly well as visible from the steep rise, but it has a similar distribution to other approaches for higher OOV scores, barring T5.
Figure 2: OOV score distribution for lipogram models. Better models exhibit left-shifted distributions.
4.4 Readability Scores
@@
To evaluate the readability of the generated text, we have utilized the TRank Readability Score Trokhymovych et al. (2024) This new multilingual model for Automatic Readability Assessment (ARA) adapts a ranking-based architecture to score individual texts, where a larger score indicates more difficult readability. Using this metric, we observed that readability is worst for the E-Delete method, followed by the synonym replacement method, with the best readability observed for the Llama3 method, as expected.
4.5 Grammatical Scores
@@
We assessed the grammatical accuracy of the generated text using Python’s LanguageTool, an open-source grammar and spell checker Mozgovoy (2011). This tool helped identify and correct grammatical errors, ensuring that the generated text adheres to standard language conventions. The score is given in terms of the average number of mistakes found in translating each source paragraph, so longer paragraphs are fairly expected to have more errors.
@@
Figure 5: Relationship between grammatical mistakes and lipogram constraint in generated texts.
The paragraph-score distributions are presented in Figure 3. Llama3 closely matches the original text in terms of grammatical correctness, followed by T5 and Llama2. The baseline E-Delete method introduces the most grammatical errors, followed by the synonym replacement method.
4.6 Ablation Study
To better understand how each part of our lipogram generation pipeline contributes to the final output, we systematically removed components and observed the effects on the first 200 paragraphs. Table 3 presents the results.
Finetuning on paraphrase data proved important to downstream performance. Finetuning keeps the generated text semantically close to the original and reducing odd, out-of-place (OOV) words. Moreover, readability worsened to -2.21, showing the importance of finetuning for maintaining both the semantic integrity and overall quality of the text. On the flip side, removing multiselection and beamsearch led only to a slight reduction in similarity to 0.68, suggesting that while these methods are beneficial, their impact is not as pronounced as finetuning. Excluding post-processing resulted in a marginal increase in the similarity score by 0.01 but caused a higher OOV percentage of 36.53% and more grammatical mistakes (2.37%).

### v2

@@ 1 Introduction
The lipogram is interesting as a highly specific literary form, one that requires the writer to omit all occurrences of a particular letter from the text. Perhaps the best known lipogram is Gadsby (Wright, 1939), which systematically excludes the letter ‘e’ throughout his novel. Crafting such a lipogram is challenging, because ‘e’ is the most common letter in English, representing roughly 12% of the letters appearing in conventional English-language text.
Lipograms provide insight into the flexibility and malleability of language under strong constraints. Modern large language models (LLMs) excel in generating text under strong stylistic constraints, such as replicating the style of Shakespeare or Dr. Seuss. This makes lipogram generation an excellent stress test for LLMs, to navigate these literary constraints without sacrificing fluency or meaning.
@@
Our particular challenge is to translate F. Scott Fitzgerald’s classic American novel “The Great Gatsby” Fitzgerald (1925) into an ‘e’-less lipogram. Gatsby was selected as our target because it is a short (47,094 words) novel familiar to a broad readership, and with the recent expiration of copyright protection now in the public domain. Finally, the name Gatsby resonates well with Gadsby, the lipogram novel which is its rough contemporary.
@@
GPT-4o During my youth, at a time of fragility, my dad imparted wisdom that I continuously mull over in thought.
@@
Our contributions in this paper include:
  • Methods for Lipogram Translation: We present a systematic study of algorithms for creating lipograms, from baseline methods such as synonym replacement to sophisticated generative models enhanced using (i) beam search, (ii) name entity analysis, (iii) training on paraphrased texts, (iv) optimization by sequence trimming and (v) multi-selection. We present an ablation study that assesses the relative effectiveness of these techniques.
  • Lipogram Evaluation: We measure the quality of generated lipogram translations across a variety of metrics, including readability, grammatical correctness, and semantic similarity to the original source text. We use this evaluation suite to compare the performance of three different LLMs, demonstrating that that Llama3 produced the most faithful as well as the most readable constrained translations.
Our project not only demonstrates the feasibility of creating high-quality lipograms using LLMs, but also provides valuable insights into the complexities of constrained text generation. We provide the first page of our lipogram translation of Gatsby for inspection in Section B of the appendix, along with the original source text.  [^1] Upon acceptance of this paper we will release our full lipogram translation and all associated data/code.
2 Previous Work
  • La Disparition (A Void) (Perec, 1969) by Georges Perec is a French novel inspired by “Gadsby" that also excludes the letter ‘e’. It was later translated into English by Gilbert Adair while maintaining this constraint. Perec followed up with his novel “Les Revenentes” Perec (1972) which uses only words with the vowel ‘e’, showcasing his exploration of constrained writing beyond lipograms.
The challenge of constrained writing has garnered significant attention in both literary circles and popular media. The Oulipo (Ouvroir de littérature potentielle) group was instrumental in exploring the potential of literary constraints. Their works exemplify the creative possibilities that constraints can unlock in literature Calvino and Weaver (1993); Motte (1986).
Recent strides in NLP have redefined constrained text generation. Roush et al. (2023) introduced token filtering for robust constraint adherence, while Kumar et al. (2021) framed it as an optimization problem. Zhang et al. (2020) proposed POINTER, an insertion-based transformer for efficient hard constraints, and Liu et al. (2021) used reinforcement learning to improve concept coverage, though it struggles with stricter constraints. Post and Vilar (2018) introduced dynamic beam allocation to enforce lexical constraints efficiently. We build on this by handling stricter character-level bans and ensuring coherent long-form generation.
3 Generation Approaches
Here we discuss our methods for translation under lipogram constraints. We will compare them against two simple baselines:
  • E-Removal: Simply removes all instances of ‘e’ from the text. This approach introduces spelling errors and reduces readability.
  • Synonym Replacement: Words were lemmatized and replaced with ‘e’-free synonyms from WordNet (Miller, 1995). If no suitable synonym existed, the original word was retained with ‘e’s removed.
  • GPT-4o Prompting: Prompts GPT-4o to rewrite text without the letter ‘e’; often fluent but may break constraints.
Fine tuning for Paraphrasing:
To enhance performance on the paraphrasing task, we fine-tuned our model using the PAWS (Zhang et al., 2019) dataset, which provides pairs of original and paraphrased texts. For fine-tuning the Llama2 model, we utilized QLoRA Dettmers et al. (2024) with 4-bit quantization and the SFTTrainer framework. For T5, the loss function was modified to include penalties for generating words with the letter ‘e’ and for maintaining high semantic similarity to the original text.
Constrained Beam Search:
To generate text that strictly avoids the letter ‘e’, we employed a constrained beam search approach during decoding. Tokens containing ‘e’ were filtered using a bad_words_ids list, and beam search was configured with 20 beams to expand the search space, and a no-repeat n-gram size of 3 was applied to avoid repetitive patterns in the output.
Multiselection:
Our model generated multiple candidate translations selecting the one with the highest cosine similarity to the original. This ensured compliance with constraints while preserving content and context, reducing randomness or irrelevance.
LLM Selection:
We experimented with both encoder-decoder (T5-small) and decoder-only (Llama2-7b-chat-hf and Llama3-8b-Instruct) models. T5 employed constrained beam search, named entity recognition (NER), and pronoun resolution, while Llama2 and Llama3 incorporated QLoRA quantization and constrained decoding.
Post Processing:
We refined the generated text to ensure clarity and readability. Overlong outputs were truncated to remove irrelevant content. Punctuation was standardized, lengthy lists were trimmed, and empty paragraphs were eliminated to maintain flow. LanguageTool corrected grammatical errors, while Proselint Richardson et al. (2024) polished the writing style and improved spacing.
Model e% CS OOV
GM
RD
Original 37.54 1.00 35.56 1.19 -3.10
E-Delete 0.00 0.61 51.32 29.26 -2.21
Synonym 0.00 0.67 40.21 24.85 -2.78
GPT-4o 5.94 0.74 40.21 1.62 -2.83
T5 0.00 0.69 35.27 10.39 -3.15
Llama2 0.00 0.65 37.41 16.40 -3.75
Llama3 0.00 0.71 36.37 1.08 -3.96
@@ 3 Generation Approaches
Model CS OOV
GM
RD
M 0.69 35.04 0.88 -4.00
M–FT 0.23 63.17 26.47 -2.21
M–(MS, BS) 0.42 35.74 0.93 -3.62
M–PS 0.43 36.53 2.37 -4.00
Table 3: Ablation studies on our Llama3-based lipogram translation model. M = Complete Model, FT = Finetuning, MS = Multiselection, BS = Beam Search, PS = Post Processing
@@ 4 Evaluation Methods
We evaluated the text on semantic similarity, readability, grammatical accuracy, and constraint adherence. All of our methods achieved 100% adherence to the ‘e’-less constraint. Table 2 highlights model performance across these metrics.
@@
Figure 1: Cosine similarity distributions for lipogram translations. Better models have right-shift distributions.
Cosine Similarity of Embeddings (CS):
@@
We used cosine similarity to measure the semantic equivalence between the original source and our generated lipogram translation. For embeddings, we utilized the the SentenceTransformer model “paraphrase-MiniLM-L6-v2" model to encode semantics (Reimers and Gurevych, 2019). Figure 1 presents a histogram of paragraph similarity over our models. GPT-4o and Llama3 generate more similar paragraphs, although there remains a sizable gap with the original Gatsby text.
Out of Vocabulary Score:
@@
Lipogram construction algorithms are naturally tempted to delete all occurrences of forbidden symbols, leaving non-vocabulary word behind (i.e. bhind). To measure this we calculated an out of vocabulary (OOV) score as the percentage of words in the generated text that were not found in standard dictionaries. From Table 2, we note that 35.56% of the words in original Gatsby qualify as OOV, largely reflecting proper nouns and obsolete terms. E-Delete inflates OOV to 51.32% due to ill-formed words, while T5 (35.27%) and Llama3 (36.37%) closely align with the original, reflecting better lexical precision.
Readability Scores (RD):
@@
To evaluate the readability of the generated text, we have utilized the TRank Readability Score Trokhymovych et al. (2024). Using this metric, we observed that readability is worst for the E-Delete method, followed by the synonym replacement method, with the best readability observed for the Llama3 method, as expected.
Grammatical Mistake (GM):
@@
We assessed the grammatical accuracy of the generated text using Python’s LanguageTool Mozgovoy (2011). The score is given in terms of the average number of mistakes found in translating each source paragraph, so longer paragraphs are fairly expected to have more errors. As presented in Table 2 (clearer in Appendix A Figure 3), Llama3 closely matches the original text in terms of grammatical correctness, followed by T5 and Llama2. The baseline E-Delete method introduces the most grammatical errors.
4.1 Ablation Study
To better understand how each part of our lipogram generation pipeline contributes to the final output, we systematically removed components and observed the effects on the first 200 paragraphs. Table 3 presents the results. Finetuning on paraphrase data proved important to downstream performance. Finetuning keeps the generated text semantically close to the original and reducing odd, out-of-place (OOV) words. Moreover, readability difficulty increased to -2.21, showing the importance of finetuning for maintaining both the semantic integrity and overall quality of the text. Removing multiselection and beamsearch also led the reduction in similarity to 0.42. Excluding post-processing resulted in a marginal increase in the similarity score by 0.01 but caused a higher OOV percentage of 36.53% and more grammatical mistakes (2.37%).